# spacing=nonfrench

### v1

@@ 2.3 Plemelj-Calderón Problem for $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$
The space $\mathcal{B}_{2,2}^{1/2}(\Gamma)$ has two remarkable properties: it is conformally invariant and its definition does not involve the distance to the boundary. As a corollary, as we have seen, Plemelj-Calderón property holds for this space for all quasicircles. For $p\neq 2$ these properties split: the conformality property remains for the spaces $\mathcal{B}_{p,p}^{1/p}(\Gamma),\,p>1$ but in its definition a distance term $d(z,\Gamma)^{p-2}$ appears. This case will be discussed in the next section, together with the general case of $s\in(0,1)$ with a distance term $d(z,\Gamma)^{(1-s)p-1}$. The other property transfers to $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$. Notice that $1-1/p=1/p^{\prime}$ where $p^{\prime}$ is the coefficient conjugate to $p$.
@@
Let $\Gamma$ be a quasicircle and $p>2$. Every $f\in\widetilde{\mathcal{B}}_{p,p}^{1-1/p}(\Gamma)$ admits a unique decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ and
@@
Before we proceed to the proof of this theorem, one must properly define what we mean by $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$. It happens that functions in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ are Hölder continuous in the closure of $\Omega_{i,e}$ with exponent $\alpha=1-2/p$. This may be seen by two different ways. Firstly, s=11/p>1/ps=1-1/p>1/p, we will prove later that functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are $\alpha$-Hölder continuous in the closure of $\Omega_{i,e}$ with $\alpha=s-1/p$. The other way to prove the same thing is to appeal to Jones’s theorem: a function in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ can be extended to a function in $W^{1,p}(\mathbb{C})$ even though clearly one cannot extend in this case by using quasiconformal reflection argument, used in the case of $p=2$ as above (see [22], also see [8]). The Morrey’s inequality states that $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$, $p>2$ (note that this inclusion relation holds in $\mathbb{R}^{n}$ for $p>n$) for $\alpha=1-2/p$, the space of $\alpha$-Hölder continuous functions in $\mathbb{C}$. Knowing this fact one may re-define the space $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$ as the space of $\Lambda^{1-2/p}$-functions on $\Gamma$ whose harmonic extensions to $\Omega_{i,e}$ belong to $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$, that is,
@@
Thanks to Jones’ theorem the proof is exactly the same as for $p=2$ because the Beurling transform is bounded on $L^{p}(\mathbb{C})$ for all $p>1$. Notice that the argument of uniqueness used in the case $p=2$ goes through here because $L^{p}(\Omega_{i})\subset L^{2}(\Omega_{i})$ for $p>2$. ∎
3 Fractional Sobolev Spaces on Quasidisks
3.1 Plemelj-Calderón Property for General $\mathcal{B}_{p,p}^{s}(\Gamma)$
As mentioned in the proceeding section, in order to deal with Plemelj-Calderón property for general $p>1,s\in(0,1)$ we need to deal with boundedness of the Beurling transform on $L^{p}(\omega,\,\mathbb{C})$ where $\omega$ is a weight. By a weight we mean a nonnegative locally integrable function on $\mathbb{C}$. Since the Beurling transform is a Calderón-Zygmund operator, we know [9] that this property holds if and only if $\omega$ belongs to the so-called class $A_{p}$ whose definition we now recall.
@@ 3.1 Plemelj-Calderón Property for General $\mathcal{B}_{p,p}^{s}(\Gamma)$
For future use let us notice that $A_{p}$ weights on $\mathbb{T}$ are defined similarly. It is known that for all $p>1$, $A_{1}\subset A_{p}\subset A_{\infty}$, $\omega\in A_{p}$ is equivalent to $\omega^{-1/(p-1)}\in A_{p^{\prime}}$ where $1/p+1/p^{\prime}=1$, and $A_{p_{1}}\subset A_{p_{2}}$ for $1<p_{1}<p_{2}$. By Hölder inequality the left part of the inequality (3.1) is always $\geq 1$: one can thus interpret this class as verifying some kind of reverse Hölder inequality. If the weight $\omega$ has $A_{p}$ then any Calderon-Zygmund operator is bounded on the weighted space $L^{p}(\omega,\mathbb{C})$ (see [9]).
@@
Notice $F_{i}^{\prime}(z)=B(\bar{\partial}f\chi_{\Omega_{e}})(z)=B(\bar{\partial}u_{e}\chi_{\Omega_{e}})(z)$, $F_{e}^{\prime}(z)=-B(\bar{\partial}f\chi_{\Omega_{i}})(z)=-B(\bar{\partial}u_{i}\chi_{\Omega_{i}})(z)$ where $u_{i,e}$ are the harmonic extensions of $f|_{\Gamma}$ to $\Omega_{i,e}$, we have that a sufficient condition on $\Gamma$ that implies $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is that the weight $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ satisfies the $A_{p}$ condition. Precisely, under this condition we have
@@
The dependence of the condition that $d(z,\Gamma)^{(1-s)p-1}$ having $A_{p}$ on the geometry of the curve has been studied by Astala [3]. In order to state his results we first need to introduce some notions.
For a compact set $E\subset\mathbb{C}$ and $0<\delta\leq 2$ let
@@
Here, $E+D(0,t)$, by the definition $\{e+z:\;e\in E,\;z\in D(0,t)\}$, is a set of points at a distance less than $t$ from $E$ which is called Minkowski sausage. We then define a kind of Minkowski content by
@@
and we may say the degree of regularity of $\Gamma$ is $h(\Gamma)$. For more information on $h(\Gamma)$, and its relation with the Hausdorff dimension $\text{dim}_{H}(\Gamma)$ of $\Gamma$, see [3].
@@
Define $\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ to be the closure of
@@
in $\mathcal{B}_{p,p}^{s}(\Gamma).$ Notice that every quasicircle is porous (see [41]). We may now state the principal theorem of this section:
@@
Let $\Gamma$ be a quasicircle and $1<p<\infty,\;0<s<1$. If the point $(h(\Gamma),s)$ locates in the shadowed region (see Figure 1); that is,
$(h(\Gamma)-1)\frac{p-1}{p}<s<\frac{p+1-h(\Gamma)}{p},$ (3.3)
then any $f\in\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ has a decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and
@@
Following the discussion above, it is clear that (3.4) holds if it holds that the weight $d(z,\Gamma)^{(1-s)p-1}\in A_{p}$. Assume first that $s>\frac{p-1}{p}$ so that (1s)p<1(1-s)p<1. By Theorem 3.4, $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow(1-s)p>h(\Gamma)-1\Leftrightarrow s<\frac{p+1-h(\Gamma)}{p}$.
Assume now $s<\frac{p-1}{p}$. Then, $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow d(z,\Gamma)^{\frac{1-(1-s)p}{p-1}}\in A_{p^{\prime}}$. Here, $\frac{1-(1-s)p}{p-1}=\frac{sp}{p-1}-1$ with $\frac{sp}{p-1}<1$. By Theorem 3.4 again we have thus that $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow s>\frac{p-1}{p}(h-1)$. The decomposition can be proved similar to that of Theorem 2.3. The case of s=11/ps=1-1/p has been addressed in the last section. ∎
We define $\mathcal{A}$ to be the admissible set involved in Theorem 3.5, that is,
@@
We end this part with a comment on the triangular domain above the shadowed region in Figure 1. Let the curve $\Gamma$ be a self-similar isotropic fractal which satisfies the open set condition. Its fractal dimension $d(\Gamma)$ equals to $h(\Gamma)$ in this case (see [32]). For instance, the Koch snowflake curve, we mentioned before, is a such quasicircle with $d(\Gamma)=\log 4/\log 3$. Denote by $N(\varepsilon)$ the minimal number of disks of radius $\varepsilon$ necessary to cover $\Gamma$. Then,
$d(\Gamma)=\lim_{\varepsilon\to 0}\frac{\log N(\varepsilon)}{\log\frac{1}{\varepsilon}}.$
Taking $\varepsilon=2^{-(n+1)}$ with sufficiently large $n$ one has the difference between the number $N(\varepsilon)$ and $2^{(n+1)d(\Gamma)}$ approaches 0 as $n$ increases to $\infty$. By dividing the domain $\Omega_{i}$ into pieces:
$\{z\in\Omega_{i}:\;d(z,\Gamma)\geq 2^{-1}\},\quad\{z\in\Omega_{i}:\;2^{-(j+1)}\leq d(z,\Gamma)<2^{-j}\},\quad j=1,2,3,\cdots$
and then computing directly one can see that the integral $\iint_{\Omega_{i}}d(z,\Gamma)^{(1-s)p-1}dxdy$ converges if and only if the series $\sum 2^{-n(p-sp+1-h(\Gamma))}$ converges, namely, $s<(p+1-h(\Gamma))/p$. Consequently, assuming $s\geq(p+1-h(\Gamma))/p$ one can see that in this case even the simplest function $u(z)=z$ is not in the space $\mathcal{B}^{s}_{p,p}(\Omega_{i})$.
3.2 Plemelj-Calderón Property for $\mathcal{B}_{p,p}^{1/p}(\Gamma)$
@@
Under the assumption of Theorem 3.6, when $s<1/p$ the only thing that we can say is that $\mathcal{B}_{p,p}^{s}(\Gamma)$ contains $\mathcal{B}_{p,p}^{1/p}(\Gamma)$ and thus also $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$.
@@
The assumption on $p$ is exactly the condition for $(h(\Gamma),p,\frac{1}{p})\in\mathcal{A}$. If it is satisfied we then have that $\omega(z)=d(z,\Gamma)^{p-2}$ is an $A_{p}$ weight, in particular, integrable locally in $\mathbb{C}$. Suppose $g\in C_{c}^{\infty}(\mathbb{C})$. Then,
@@
For a general $f\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$, we will use the argument on the boundary values of $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ in the first paragraph of section 2.1. Let $u$ be the harmonic extension to $\Omega_{i}$ of $f$. Noting that $\omega\in A_{p}$ we see that $u$ has an extension $U\in W^{1,p}(\omega,\;\mathbb{C})$ to the whole plane (see [8]), and there exists a sequence $U_{n}$ in $C_{c}^{\infty}(\mathbb{C})$ such that (see Lemma 3.13 and also [8])
@@
We already know that this decomposition is unique if $p=2$. Since the spaces $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ increase with $p$ we still have uniqueness for $h(\Gamma)<p\leq 2$.

### v2

@@ 2.4 Plemelj-Calderón problem for $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$
The space $\mathcal{B}_{2,2}^{1/2}(\Gamma)$ has two remarkable properties: it is conformally invariant and its definition does not involve the distance to the boundary. As we have seen, Plemelj-Calderón property holds for this space for all quasicircles. For $p\neq 2$ these properties split: the conformality property remains for the spaces $\mathcal{B}_{p,p}^{1/p}(\Gamma),\,p>1$, but in its definition a distance term $d(z,\Gamma)^{p-2}$ appears. This case will be discussed in the next subsection, together with the general case of $s\in(0,1)$ with a distance term $d(z,\Gamma)^{(1-s)p-1}$. The other property transfers to $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$. Notice that $1-1/p=1/p^{\prime}$ where $p^{\prime}$ is the coefficient conjugate to $p$.
@@
Let $\Gamma$ be a quasicircle and $p>2$. Each $f\in\widetilde{\mathcal{B}}_{p,p}^{1-1/p}(\Gamma)$ admits a unique decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ and
@@
Before we proceed to the proof of this theorem, one must properly define what we mean by $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$. It happens that functions in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ are Hölder continuous in the closure of $\Omega_{i,e}$ with exponent $\alpha=1-2/p$. This may be seen by two different ways. Firstly, s=11/p>1/ps=1-1/p>1/p, we will prove later that the functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are $\alpha$-Hölder continuous in the closure of $\Omega_{i,e}$ with $\alpha=s-1/p$. The other way to prove the same thing is to appeal to Jones’ theorem: a function in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ can be extended to a function in $W^{1,p}(\mathbb{C})$ even though clearly one cannot extend in this case by using quasiconformal reflection argument, used in the case of $p=2$ as above (see [22], also see [8]). The Morrey’s inequality states that $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$, $p>2$ (note that this inclusion relation holds in $\mathbb{R}^{n}$ for $p>n$) for $\alpha=1-2/p$, the space of $\alpha$-Hölder continuous functions in $\mathbb{C}$. Knowing this fact, one may re-define the space $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$ as the space of $\Lambda^{1-2/p}$-functions on $\Gamma$ whose harmonic extensions to $\Omega_{i,e}$ belong to $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$, that is,
@@
Thanks to Jones’ theorem, the proof is exactly the same as that for $p=2$ because the Beurling transform is bounded on $L^{p}(\mathbb{C})$ for all $p>1$. Notice that the argument of uniqueness used in the case $p=2$ goes through here because $L^{p}(\Omega_{i})\subset L^{2}(\Omega_{i})$ for $p>2$, and $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{e})\subset\mathcal{B}_{2,2}^{1/2}(\Omega_{e})$ by the argument in Appendix. ∎
2.5 Plemelj-Calderón problem for general $\mathcal{B}_{p,p}^{s}(\Gamma)$
In order to deal with Plemelj-Calderón property for general $p>1,s\in(0,1)$ we need to consider the boundedness of the Beurling transform on $L^{p}(\omega,\,\mathbb{C})$ where $\omega$ is a weight. By weight we mean a nonnegative locally integrable function in $\mathbb{C}$. Since the Beurling transform is a Calderón-Zygmund operator, we know [9] that this property holds if and only if $\omega$ belongs to the so-called class $A_{p}$ whose definition we now recall.
@@ 2.5 Plemelj-Calderón problem for general $\mathcal{B}_{p,p}^{s}(\Gamma)$
For future use, let us note that the $A_{p}$-weights on $\mathbb{T}$ are defined similarly. It is known that for all $p>1$, $A_{1}\subset A_{p}\subset A_{\infty}$, $\omega\in A_{p}$ is equivalent to $\omega^{-1/(p-1)}\in A_{p^{\prime}}$ where $1/p+1/p^{\prime}=1$, and $A_{p_{1}}\subset A_{p_{2}}$ for $1<p_{1}<p_{2}$. By Hölder inequality, the left part of the inequality (2.15) is always not less than $1$: thus one can interpret this class as verifying some kind of reverse Hölder inequality. If the weight $\omega$ has $A_{p}$ then any Calderon-Zygmund operator is bounded on the weighted space $L^{p}(\omega,\mathbb{C})$ (see [9]).
@@
Notice
$F_{i}^{\prime}(z)=B(\bar{\partial}f\chi_{\Omega_{e}})(z)=B(\bar{\partial}u_{e}\chi_{\Omega_{e}})(z)$
and
$F_{e}^{\prime}(z)=-B(\bar{\partial}f\chi_{\Omega_{i}})(z)=-B(\bar{\partial}u_{i}\chi_{\Omega_{i}})(z)$
by Lemma 2.2, where $u_{i,e}$ are the harmonic extensions of $f|_{\Gamma}$ to $\Omega_{i,e}$, we see that a sufficient condition on $\Gamma$ which implies $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is that the weight $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ satisfies the $A_{p}$ condition. Precisely, under this condition we have the following.
@@
The dependence of the condition $d(z,\Gamma)^{(1-s)p-1}$ having $A_{p}$ on the geometry of the curve has been studied by Astala [3]. In order to state his results, we first need to introduce some notions.
For a compact set $E\subset\mathbb{C}$ and $0<\delta\leq 2$, let
@@
Here, $E+D(0,t)$, by the definition $\{e+z:\;e\in E,\;z\in D(0,t)\}$, is a set of points at a distance less than $t$ from $E$, which is called Minkowski sausage. (Here and in the sequel we write $|E+D(0,t)|$ for the Lebesgue measure of the set $E+D(0,t)$.) We then define a kind of Minkowski content by
@@
We mention here that in the definition of $h_{\delta}(E)$, $M_{\delta}(E,t)$ is sometimes replaced in the literature by
$H_{\delta}(E,t):=\inf\{nt^{\delta}:\;E\subset\cup_{i=1}^{n}D(z_{i},t),\;n\in\mathbb{N}\}.$
This replacement does not make any difference since there are universal constants $c_{1}$ and $c_{2}$ such that (see [36])
$c_{1}M_{\delta}(E,t)\leq H_{\delta}(E,t)\leq c_{2}M_{\delta}(E,t).$
@@
and we may say that the degree of regularity of $\Gamma$ is $h(\Gamma)$. For more information on $h(\Gamma)$, and its relation with the Hausdorff dimension $\text{dim}_{H}(\Gamma)$ of $\Gamma$; see [3].
@@
Define $\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ as the closure of
@@
in $\mathcal{B}_{p,p}^{s}(\Gamma).$ Notice that every quasicircle is porous (see [47]). We may now state the principal theorem of this subsection.
@@
Let $\Gamma$ be a quasicircle and $1<p<\infty,\;0<s<1$. If the point $(h(\Gamma),s)$ is located in the shadowed region (see Figure 1); that is,
$(h(\Gamma)-1)\frac{p-1}{p}<s<\frac{p+1-h(\Gamma)}{p}$ (2.17)
(that would be simplified to $0<s<1$ if $h(\Gamma)=1$, i.e. $\Gamma$ is a chord-arc curve ), then any $f\in\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ has a decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and
@@
Following the discussion above, clearly (2.18) holds if $d(z,\Gamma)^{(1-s)p-1}\in A_{p}$ holds. Assume first that $s>\frac{p-1}{p}$ so that (1s)p<1(1-s)p<1. By Theorem 2.8, $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow(1-s)p>h(\Gamma)-1\Leftrightarrow s<\frac{p+1-h(\Gamma)}{p}$.
Assume now $s<\frac{p-1}{p}$. Then $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow d(z,\Gamma)^{\frac{1-(1-s)p}{p-1}}\in A_{p^{\prime}}$. Here, $\frac{1-(1-s)p}{p-1}=\frac{sp}{p-1}-1$ with $\frac{sp}{p-1}<1$. By Theorem 2.8 again we have thus $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow s>\frac{p-1}{p}(h-1)$. The decomposition can be shown to be similar to that of Theorem 2.3. The case of s=11/ps=1-1/p has been addressed in the last subsection. ∎
We define $\mathcal{A}$ as the admissible set involved in Theorem 2.9, that is,
@@
Remark 2.10.
We end this subsection with a comment on the triangular domain above the shadowed region in Figure 1. Divide the domain $\Omega_{i}$ into pieces:
$E_{0}$ $=\{z\in\Omega_{i}:\;d(z,\Gamma)\geq 2^{-1}\},$
$E_{n}$ $=\{z\in\Omega_{i}:\;2^{-(n+1)}\leq d(z,\Gamma)<2^{-n}\},\quad n=1,2,3,\cdots$
If $t<R$ it follows from basic covering theorems that there exists a constant $C_{1}$ such that (see p.474 in [3])
$t^{h(\Gamma)-2}|\{z\in D(z_{0},R):\;d(z,\Gamma)<t\}|\leq C_{1}R^{h(\Gamma)}$
where $D(z_{0},R)$ stands for the open disk centered at $z_{0}$ and of radius $R$. From this we see $|E_{n}|\leq C_{2}2^{-n(2-h(\Gamma))}$. Then, by direct computation one can see that the integral $\iint_{\Omega_{i}}d(z,\Gamma)^{(1-s)p-1}dxdy$ converges if and only if the series $\sum 2^{-n(p-sp+1-h(\Gamma))}$ converges, namely, $s<(p+1-h(\Gamma))/p$. Consequently, assuming $s\geq(p+1-h(\Gamma))/p$ one can see that even the simplest function $u(z)=z$ is not in the space $\mathcal{B}^{s}_{p,p}(\Omega_{i})$.
2.6 Plemelj-Calderón problem for $\mathcal{B}_{p,p}^{1/p}(\Gamma)$
@@
Under the assumption of Theorem 2.11, when $s<1/p$ the only thing we can say is that $\mathcal{B}_{p,p}^{s}(\Gamma)$ contains $\mathcal{B}_{p,p}^{1/p}(\Gamma)$ and thus also $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$.
@@
The assumption on $p$ is exactly the condition for $(h(\Gamma),p,\frac{1}{p})\in\mathcal{A}$. If it is satisfied we then have that $\omega(z)=d(z,\Gamma)^{p-2}$ is an $A_{p}$-weight, in particular, integrable locally in $\mathbb{C}$. Suppose $g\in C_{c}^{\infty}(\mathbb{C})$. Then,
@@
For a general $f\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$, we will use the argument on the boundary values of $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ in the first paragraph of subsection 2.1. Let $u$ be the harmonic extension to $\Omega_{i}$ of $f$. Noting that $\omega\in A_{p}$ we see that $u$ has an extension $U\in W^{1,p}(\omega,\;\mathbb{C})$ to the whole plane (see [8]), and there exists a sequence $U_{n}$ in $C_{c}^{\infty}(\mathbb{C})$ such that (see Lemma 3.3 and also [8])
@@
We already know that this decomposition is unique if $p=2$. Since the spaces $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ increase with $p$, we still have uniqueness for $h(\Gamma)<p\leq 2$.
3 $A_{p}$-Weighted Sobolev spaces on quasidisks